OpenStudy (anonymous):

3 c 6 fl oz + 5 c 6 fl oz

OpenStudy (anonymous):

these are answers 8 c 4 fl oz 9 c 10 fl oz 9 c 4 fl oz

OpenStudy (dmezzullo):

3c +5 c 6 +6 8c 12fl

OpenStudy (dmezzullo):

Those make no sense it is 8 and 12

OpenStudy (anonymous):

what

OpenStudy (dmezzullo):

It is 8 c 12 fl oz

OpenStudy (anonymous):

that's not m answer i only have 8 c 4 fl oz 9 c 10 fl oz 9 c 4 fl oz

OpenStudy (dmezzullo):

Well those aint right

OpenStudy (anonymous):

yes they are

OpenStudy (dmezzullo):

5+3=8

jimthompson5910 (jim_thompson5910):

use the idea that 1 cup = 8 fluid ounces

OpenStudy (dmezzullo):

Oh nvm.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so the answer is 8 c 4 fl oz

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

oh can u help me

jimthompson5910 (jim_thompson5910):

dmezzullo has the right idea

jimthompson5910 (jim_thompson5910):

the system is just picky and wants it in a certain specific form

OpenStudy (anonymous):

yeah i know and that is what is tricky

jimthompson5910 (jim_thompson5910):

8 c + 12 fl oz = 8 c + 8 fl oz + 4 fl oz 8 c + 12 fl oz = 8 c + 1 c + 4 fl oz 8 c + 12 fl oz = 9 c + 4 fl oz

jimthompson5910 (jim_thompson5910):

basically those 8 fl oz carry over and fill one extra cup

OpenStudy (anonymous):

but these r the answers 8 c 4 fl oz 9 c 10 fl oz 9 c 4 fl oz

jimthompson5910 (jim_thompson5910):

yes what I wrote is listed

OpenStudy (anonymous):

i don't see it lol

jimthompson5910 (jim_thompson5910):

9 c + 4 fl oz = ???

jimthompson5910 (jim_thompson5910):

it's pretty much there